# Sparse Extended Mean-Variance-CVaR Portfolios with Short-selling

### v1

@@ 3.2 Analysis of Algorithm 2
Here, we establish that our proposed customized penalty decomposition Algorithm 2 obtains a local minimum of the original nondifferentiable nonconvex problem ($P$). Recall that Robinson’s constraint qualification conditions for a local minimizer $(x^{*},\gamma^{*})$ of ($P$) is the existence of an index set $\mathcal{L}$ such that $|\mathcal{L}|=k$ and $x^{*}_{\mathcal{L}^{c}}=0$ such that the following holds [11]:
@@
$\left\{\begin{bmatrix}-d-v\\ d-\bar{v}\\ -d-\hat{v}\\ d-\tilde{v}\\ \textbf{e}^{T}d\\ d_{\mathcal{L}^{c}}\end{bmatrix}\ \big{|}\ \left\{\begin{array}[]{ll}d\in\mathbb{R}^{n}\\ v\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad v_{i}\leq 0;\,\forall i\in\{i\,|\,x_{i}=L_{i}\}\\ \bar{v}\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad\bar{v}_{i}\leq 0;\,\forall i\in\{i\,|\,x_{i}=U_{i}\}\\ \hat{v}=[\hat{v}_{\Gamma^{+}};0]\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad\hat{v}_{i}\leq 0;\,\forall i\in\Gamma^{+}\cap\mathcal{L}^{c}\\ \tilde{v}=[\tilde{v}_{\Gamma^{-}};0]\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad\hat{v}_{i}\leq 0;\,\forall i\in\Gamma^{-}\cap\mathcal{L}^{c}\end{array}\right.\right\}=\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}^{|\mathcal{L}^{c}|}$ (26)
@@
Through the linearity of constraints in ($P$) except the cardinality constraint, it is easy to show that Robinson’s conditions above always hold for an arbitrary $x^{*}$. Under these Robinson’s conditions, the KKT conditions for a local minimizer $(x^{*},\gamma^{*})$ of ($P$) are the existence of Lagrangian multipliers $(\bar{t},\bar{T}_{1},\bar{T}_{2},\bar{T}_{5},\bar{T}_{6})$ with $\bar{t}\in\mathbb{R},\bar{T}_{i}\in\mathbb{R}^{n};\forall i\in\{1,2,5,6\},$ and $\bar{\Omega}\in\mathbb{R}^{n}$ with $\mathcal{L}\subseteq[n]$ such that $|\mathcal{L}|=k$ and the following holds:
@@
$\left\{\begin{array}[]{ll}2\lambda_{1}Ax^{*}-\lambda_{2}\Psi+\bar{t}\textbf{e}-\bar{T}_{5}+\bar{T}_{6}+\bar{\Omega}+\lambda_{2}\delta\partial\left(\|\cdot\|_{1})\right|_{x^{*}-\phi}-\bar{T}_{1}+\bar{T}_{2}+\dots\vspace{0.2cm}\\ \qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \quad\quad+\dots-\frac{\lambda_{3}}{m(1-\beta)}\sum_{j=1}^{m}d_{j}\partial\left(.)^{+}\right|_{-d_{j}^{T}x^{*}-\gamma^{*}}=0,\vspace{.1cm}\\ \lambda_{3}-\frac{1}{m(1-\beta)}\sum_{j=1}^{m}\partial\left(.)^{+}\right|_{-d_{j}^{T}x^{*}-\gamma^{*}}=0,\vspace{.2cm}\\ \textbf{e}^{T}x^{*}=1,\quad 0\leq\bar{T}_{1}\perp(x^{*}-L)\geq 0,\quad 0\leq\bar{T}_{2}\perp(U-x^{*})\geq 0,\quad x^{*}_{{\mathcal{L}}^{c}}=0,\quad\bar{\Omega}_{\mathcal{L}}=0,\vspace{0.1cm}\\ 0\leq\bar{T}_{5}\perp[(x^{*})^{+}_{\Gamma^{+}};0]\geq 0,\quad 0\leq\bar{T}_{6}\perp-[(x^{*})^{-}_{\Gamma^{-}};0]\geq 0,\end{array}\right.$ (28)
@@
Suppose that $\lambda_{1}$ and $\lambda_{2}\geq 0,$ with $\lambda_{3}=1-\lambda_{1}-\lambda_{2}>0,\beta\in(0,1),\delta\geq 0$, $\eta,L,U\in\mathbb{R}^{n}$ with $L<U$ element-wise, $d_{j}\in\mathbb{R}^{n}$ for each $j\in[m],A\succeq 0,$ and $\rho\geq 1$. Let $\left\{\left(x^{(j)},y^{(j)},z^{(j)},w^{(j)},\gamma^{(j)}\right)\right\}$ be a sequence generated by Algorithm 2 for solving ($P$). Then, the following hold:
@@
    Suppose that Robinson’s condition given in (26) holds at $(x^{*},\gamma^{*})$ with the index subset $\mathcal{L}$ indicated above. Then, $(x^{*},\gamma^{*})$ is a local minimizer of satisfying ($P$).
@@
$\lambda_{1}{x^{(j)}}^{T}Ax^{(j)}-\lambda_{2}\Psi^{T}x^{(j)}+\lambda_{2}\delta\|z^{(j)}-\phi\|_{1}+\lambda_{3}(\gamma+\frac{1}{m(1-\beta)}\sum_{j=1}^{m}(-d_{j}^{T}w^{(j)}-\gamma)^{+})\geq\hat{C}>-\infty.$
@@
$\left\{\begin{array}[]{ll}2\lambda_{1}Ax^{(j_{\ell})}-\lambda_{2}\Psi+2\rho^{(j_{\ell})}\underbrace{(3x^{(j_{\ell})}-y^{(j_{\ell})}-z^{(j_{\ell})}-w^{(j_{\ell})})}_{=(x^{(j_{\ell})}-y^{(j_{\ell})})+(x^{(j_{\ell})}-z^{(j_{\ell})})+(x^{(j_{\ell})}-w^{(j_{\ell})})}+t^{(j_{\ell})}\textbf{e}=0,\vspace{0.25cm}\\ y^{(j_{\ell})}_{\mathcal{L}}=([x^{+}_{\Gamma^{+}};0]-[x^{-}_{\Gamma^{-}};0]+[x_{\Gamma^{0}};0])^{(j_{\ell})}_{\mathcal{L}},\vspace{0.25cm}\\ 2\rho^{(j_{\ell})}(z^{(j_{\ell})}-x^{(j_{\ell})})+\lambda_{2}\delta\partial\left(\|\cdot\|_{1})\right|_{z^{(j_{\ell})}-\phi}-T^{(j_{\ell})}_{1}+T^{(j_{\ell})}_{2}=0,\vspace{0.25cm}\\ \frac{-\lambda_{3}}{m(1-\beta)}\sum_{j=1}^{m}d_{j}\partial\left(.)^{+}\right|_{-d_{j}^{T}w^{(j_{\ell})}-\gamma^{(j_{\ell})}}+2\rho(w^{(j_{\ell})}-x^{(j_{\ell})})-T^{(j_{\ell})}_{3}+T^{(j_{\ell})}_{4}=0,\vspace{.25cm}\\ \lambda_{3}-\frac{1}{m(1-\beta)}\sum_{j=1}^{m}\partial\left(.)^{+}\right|_{-d_{j}^{T}w^{(j_{\ell})}-\gamma^{(j_{\ell})}}=0,\vspace{0.25cm}\\ \textbf{e}^{T}x^{(j_{\ell})}=1,\quad 0\leq T^{(j_{\ell})}_{1}\perp(z^{(j_{\ell})}-L)\geq 0,\quad 0\leq T^{(j_{\ell})}_{2}\perp(U-z^{(j_{\ell})})\geq 0,\quad 0\leq T^{(j_{\ell})}_{3}\perp(w^{(j_{\ell})}-L)\geq 0,\vspace{0.25cm}\\ \mbox{and}\quad 0\leq T^{(j_{\ell})}_{4}\perp(U-w^{(j_{\ell})})\geq 0.\end{array}\right.$ (29)
Since for each $j_{\ell}$ we have $y^{(j_{\ell})}_{i}\geq 0;\forall i\in\Gamma^{+}$ and $y^{(j_{\ell})}_{i}\leq 0;\forall i\in\Gamma^{-}$, from the part (i), we can conclude $x^{(j_{\ell})}_{i}\geq 0;\forall i\in\Gamma^{+}$ and $x^{(j_{\ell})}_{i}\leq 0;\forall i\in\Gamma^{-}$. Further, $\textbf{e}^{T}x^{(j_{\ell})}=1$ and $x^{(j_{\ell})}-L\geq 0$ with $U-x^{(j_{\ell})}\geq 0$. Next, by injecting the second, third, and fourth equations of (29) into its first one, we get the following:
@@
Therefore, we (30) reduces to:
$2\lambda_{1}Ax^{(j_{\ell})}-\lambda_{2}\Psi+t^{(j_{\ell})}\textbf{e}$displaystyle-- $\underbrace{\begin{bmatrix}2\rho^{(j_{\ell})}{[x^{-}_{\Gamma^{+}};0]}^{(j_{\ell})}_{\mathcal{L}}\\ 0\end{bmatrix}}_{:=T_{5}^{(j_{\ell})}}+\underbrace{\begin{bmatrix}2\rho^{(j_{\ell})}{[x^{+}_{\Gamma^{-}};0]}^{(j_{\ell})}_{\mathcal{L}}\\ 0\end{bmatrix}}_{:=T_{6}^{(j_{\ell})}}+\underbrace{\begin{bmatrix}0\\ 2\rho^{(j_{\ell})}(x^{(j_{\ell})})_{\mathcal{L}^{c}}\end{bmatrix}}_{:=\Omega^{(j_{\ell})}}$ (31)
@@
We next show that $\{(t^{(j_{\ell})},T_{1}^{(j_{\ell})}+T_{3}^{(j_{\ell})},T_{2}^{(j_{\ell})}+T_{4}^{(j_{\ell})},T_{5}^{(j_{\ell})},T_{6}^{(j_{\ell})},\Omega^{(j_{\ell})}\}$ is bounded under Robinson’s condition on $(x^{*},\gamma^{*})$. Suppose not and consider the following normalized sequence:
@@
Since this normalized sequence is bounded, it has a convergent subsequence, without loss of generality itself, $(\widetilde{t}^{(j_{\ell})},\widetilde{T}_{1}^{(j_{\ell})}+\widetilde{T}_{3}^{(j_{\ell})},\widetilde{T}_{2}^{(j_{\ell})}+\widetilde{T}_{4}^{(j_{\ell})},\widetilde{T}_{5}^{(j_{\ell})},\widetilde{T}_{6}^{(j_{\ell})},\widetilde{\Omega}^{(j_{\ell})})$ whose limit is given by $(\widetilde{t}^{*},\widetilde{T}_{1}^{*}+\widetilde{T}_{3}^{*},\widetilde{T}_{2}^{*}+\widetilde{T}_{4}^{*},\widetilde{T}_{5}^{*},\widetilde{T}_{6}^{*},\widetilde{\Omega}^{*})$ such that $\|(\widetilde{T}_{1}^{*}+\widetilde{T}_{3}^{*},\widetilde{T}_{2}^{*}+\widetilde{T}_{4}^{*},\widetilde{T}_{5}^{*},\widetilde{T}_{6}^{*},\widetilde{\Omega}^{*})\|_{2}=1$. After dividing both sides of (31) by $\|(t^{(j_{\ell})},T_{1}^{(j_{\ell})}+T_{3}^{(j_{\ell})},T_{2}^{(j_{\ell})}+T_{4}^{(j_{\ell})},T_{5}^{(j_{\ell})},T_{6}^{(j_{\ell})},\Omega^{(j_{\ell})})\|_{2}$ and then passing the limit $j_{\ell}\to\infty$, in virtue of the part (i) and boundedness of the remaining terms in (31), continuity of .$\|.\|_{1}$ and $\max(x,0)$, and boundedness of their subgradients, we get:
$\widetilde{t}^{*}\textbf{e}-\widetilde{T}_{5}^{*}+\widetilde{T}_{6}^{*}+\widetilde{\Omega}^{*}-\widetilde{T}_{1}^{*}+\widetilde{T}_{2}^{*}-\widetilde{T}_{3}^{*}+\widetilde{T}_{4}^{*}=0,$ (32)
where $\widetilde{T}^{*}_{i}\geq 0$ with $(\widetilde{T}^{*}_{i})_{\mathcal{L}^{c}}=0;$ for $i=1,2$ and $\widetilde{T}_{i}^{*}\geq 0,$ for $i=1,2,3,4$ and $\widetilde{\Omega}^{*}_{\mathcal{L}}=0$. In light of Robinson’s condition (26), there exists $(d,v,\bar{v},\hat{v},\tilde{v})$ such that $\textbf{e}^{T}d=-\widetilde{t}^{*},-d-\hat{v}=-\widetilde{T}_{5}^{*},d-\tilde{v}=-T_{6}^{*},-d-v=-\widetilde{T}^{*}_{1}-\widetilde{T}^{*}_{3},d-\bar{v}=-\widetilde{T}^{*}_{2}-\widetilde{T}^{*}_{4},$ and $d_{\mathcal{L}^{c}}=-\widetilde{\Omega}^{*}$. Since $d_{\mathcal{L}^{c}}=-\widetilde{\Omega}^{*}_{\mathcal{L}^{c}}$ and $\widetilde{\Omega}^{*}_{\mathcal{L}}=0$, we see that $d^{T}\widetilde{\Omega}^{*}=-\|\widetilde{\Omega}^{*}\|^{2}_{2}$. By multiplying (32) and using the equations above, we get the following:
$-(\widetilde{t}^{*})^{2}-(\widetilde{T}^{*}_{5}-\hat{v})^{T}\widetilde{T}^{*}_{5}+(\tilde{v}-\widetilde{T}^{*}_{6})^{T}\widetilde{T}^{*}_{6}-(\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3}-v)^{T}(\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3})+(\bar{v}-\widetilde{T}^{*}_{2}-\widetilde{T}^{*}_{4})^{T}(\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4})=0,$
@@
$\|((\widetilde{t}^{*})^{2},\widetilde{T}^{*}_{5},\widetilde{T}^{*}_{6},\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3},\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4})\|_{2}^{2}=\underbrace{\hat{v}^{T}\widetilde{T}_{5}^{*}}_{\leq 0}+\underbrace{\tilde{v}^{T}\widetilde{T}^{*}_{6}}_{\leq 0}+\underbrace{v^{T}(\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3})}_{\leq 0}+\underbrace{\bar{v}^{T}(\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4})}_{\leq 0};$
implying that $\|((\widetilde{t}^{*})^{2},\widetilde{T}^{*}_{5},\widetilde{T}^{*}_{6},\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3},\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4})\|_{2}^{2}=0$. This is a contradiction! Consequently, the sequence $\{(t^{(j_{\ell})},T_{1}^{(j_{\ell})}+T_{3}^{(j_{\ell})},T_{2}^{(j_{\ell})}+T_{4}^{(j_{\ell})},T_{5}^{(j_{\ell})},T_{6}^{(j_{\ell})},\Omega^{(j_{\ell})}\}$ is bounded and has a convergent subsequence. Let $\{(\bar{t},\bar{T}_{1},\bar{T}_{2},\bar{T}_{5},\bar{T}_{6},\bar{\Omega})$ be the limit point of this sequence, through passing limit in (31) and applying the results of the part (i), one can see the KKT conditions (29) are satisfied.
So far, we have proved that all the KKT conditions in (29) hold at $(x^{*},\gamma^{*})$ except its second equation. To show this, recall the fourth equation in (30):
$\lambda_{3}-\frac{1}{m(1-\beta)}\sum_{j=1}^{m}\partial\left(.)^{+}\right|_{-d_{j}^{T}w^{(j_{\ell})}-\gamma^{(j_{\ell})}}=0.$
By taking limit as $j_{\ell}\to\infty$, boundedness of $\gamma^{(j_{\ell})}$ shown in (3.1), continuity of $\max(x,0)$, and considering part (i), we see that
$\lambda_{3}-\frac{1}{m(1-\beta)}\sum_{j=1}^{m}\partial\left(.)^{+}\right|_{-d_{j}^{T}x^{*}-\gamma^{*}}=0.$
Therefore, we showed all the KKT conditions hold at $(x^{*},\gamma^{*})$. Together with the linearity of constraints in ($P$) except the cardinality one, we conclude that Algorithm 2 obtains a local minimizer of this problem under the given assumptions. ∎

### v2

@@ 3.2 Analysis of Algorithm 2
Here, we establish that our proposed customized penalty decomposition Algorithm 2 obtains a Lu–Zhang minimum of the original nondifferentiable nonconvex problem ($P$). For our analysis, we need Robinson’s constraint qualification. Specifically, consider the general problem
$\min_{x\in\mathcal{X}}\,F(x)\qquad\text{subject to}\qquad G(x)\leq 0,\quad H(x)=0,\quad\text{and}\quad\|x\|_{0}\leq k,$ (12)
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is a closed convex set, and $G:\mathbb{R}^{n}\to\mathbb{R}^{m}$, $H:\mathbb{R}^{n}\to\mathbb{R}^{p}$ are continuously differentiable. Let $x^{*}$ be a feasible point. Choose an index set $\mathcal{L}\subseteq\{1,\dots,n\}$ with $|\mathcal{L}|=k$ such that $x^{*}_{j}=0$ for all $j\notin\mathcal{L}$. Define the active inequality set
$A(x^{*}):=\{\,i\in\{1,\dots,m\}:G_{i}(x^{*})=0\,\}.$
With $\overline{\mathcal{L}}=\{1,\dots,n\}\setminus\mathcal{L}$, Robinson’s constraint qualification at $x^{*}$ is the surjectivity condition below:
@@
$\left\{\begin{bmatrix}G^{\prime}(x^{*})d-v\\[4.0pt] H^{\prime}(x^{*})d\\[4.0pt] (I_{\overline{\mathcal{L}}})^{\!T}d\end{bmatrix}:\;d\in T_{\mathcal{X}}(x^{*}),\;v\in\mathbb{R}^{m},\;v_{i}\leq 0\ \forall i\in A(x^{*})\right\}=\mathbb{R}^{m}\times\mathbb{R}^{p}\times\mathbb{R}^{\,n-|\mathcal{L}|}.$
Here $G^{\prime}(x^{*})$ and $H^{\prime}(x^{*})$ denote the Jacobians of $G$ and $H$ at $x^{*}$, and $I_{\overline{\mathcal{L}}}$ is the coordinate selection matrix extracting the zero coordinates. Further, for a closed convex set $\mathcal{X}$ and $x^{*}\in\mathcal{X}$, the tangent cone is
$T_{\mathcal{X}}(x^{*})=\operatorname{cl}\Big\{\,d:\ \exists\,t_{k}\downarrow 0,\ d_{k}\to d\ \text{with}\ x^{*}+t_{k}d_{k}\in\mathcal{X}\,\Big\}.$
Because all constraints in ($P$) except the sparsity constraint are affine, the linearized operator has full row span. Hence, condition (18) holds automatically at any feasible point $x^{*}$. Moreover, since $\mathcal{X}=\mathbb{R}^{n}$, we have $T_{\mathcal{X}}(x^{*})=\mathbb{R}^{n}$. Hence, Robinson’s constraint qualification for a Lu–Zhang minimizer $(x^{*},\gamma^{*})$ of ($P$) requires the existence of an index set $\mathcal{L}\subseteq\{1,\dots,n\}$ with $|\mathcal{L}|=k$ and $x^{*}_{\mathcal{L}^{c}}=0$, ensuring that the following condition is satisfied [12]:
$\left\{\begin{bmatrix}-d-v\\ d-\bar{v}\\ -d-\hat{v}\\ d-\tilde{v}\\ \textbf{e}^{T}d\\ d_{\mathcal{L}^{c}}\end{bmatrix}\ \big|\ \left\{\begin{array}[]{ll}d\in\mathbb{R}^{n}\\ v\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad v_{i}\leq 0;\,\forall i\in\{i\,|\,x^{*}_{i}=L_{i}\}\\ \bar{v}\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad\bar{v}_{i}\leq 0;\,\forall i\in\{i\,|\,x^{*}_{i}=U_{i}\}\\ \hat{v}=[\hat{v}_{\Gamma^{+}};0]\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad\hat{v}_{i}\leq 0;\,\forall i\in\Gamma^{+}\cap\mathcal{L}^{c}\\ \tilde{v}=[\tilde{v}_{\Gamma^{-}};0]\in\mathbb{R}^{n}\quad\mbox{s.t.}\quad\tilde{v}_{i}\leq 0;\,\forall i\in\Gamma^{-}\cap\mathcal{L}^{c}\end{array}\right.\right\}=\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}\times\mathbb{R}\times\mathbb{R}^{|\mathcal{L}^{c}|},$ (18)
with $n_{1}=|\Gamma^{+}\cap\mathcal{L}^{c}|$ and $n_{2}={|\Gamma^{-}\cap\mathcal{L}^{c}|}$.
Note that although $x^{*}$ does not appear explicitly in (18), since all constraints except the cardinality constraint are affine, the condition is entirely defined at $x^{*}$. Specifically, (i) the Jacobians are evaluated at $x^{*}$, (ii) the active set depends on which inequalities are tight at $x^{*}$, and (iii) the support set $\mathcal{L}$ is chosen to match the nonzero elements of $x^{*}$. Hence, Robinson’s constraint qualification is a local condition at the reference point.
@@
Under these Robinson’s conditions, the KKT conditions for a Lu–Zhang minimizer $(x^{*},\gamma^{*})$ of ($P$) are the existence of Lagrangian multipliers $(\bar{t},\bar{T}_{1},\bar{T}_{2},\bar{T}_{5},\bar{T}_{6})$ with $\bar{t}\in\mathbb{R},\bar{T}_{i}\in\mathbb{R}^{n};\forall i\in\{1,2,5,6\},$ and $\bar{\Omega}\in\mathbb{R}^{n}$ with $\mathcal{L}\subseteq[n]$ such that $|\mathcal{L}|=k$ and the following holds:
@@
$\left\{\begin{array}[]{ll}2\lambda_{1}Ax^{*}-\lambda_{2}\Psi+\bar{t}\textbf{e}-\bar{T}_{5}+\bar{T}_{6}+\bar{\Omega}+\lambda_{2}\delta\partial\left(\|\cdot\|_{1})\right|_{x^{*}-\phi}-\bar{T}_{1}+\bar{T}_{2}-\vskip 5.69046pt\\ \qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \quad\quad\quad\frac{\lambda_{3}}{m(1-\beta)}\sum_{j=1}^{m}d_{j}\partial\left(.)^{+}\right|_{-d_{j}^{T}x^{*}-\gamma^{*}}\ni 0,\vskip 2.84544pt\\ \lambda_{3}-\frac{1}{m(1-\beta)}\sum_{j=1}^{m}\partial\left(.)^{+}\right|_{-d_{j}^{T}x^{*}-\gamma^{*}}\ni 0,\vskip 5.69046pt\\ \textbf{e}^{T}x^{*}=1,\quad 0\leq\bar{T}_{1}\perp(x^{*}-L)\geq 0,\quad 0\leq\bar{T}_{2}\perp(U-x^{*})\geq 0,\quad x^{*}_{{\mathcal{L}}^{c}}=0,\quad\bar{\Omega}_{\mathcal{L}}=0,\vskip 2.84544pt\\ 0\leq\bar{T}_{5}\perp[(x^{*})^{+}_{\Gamma^{+}};0]\geq 0,\quad 0\leq\bar{T}_{6}\perp-[(x^{*})^{-}_{\Gamma^{-}};0]\geq 0,\end{array}\right$. (20)
@@
Suppose that $\lambda_{1}$ and $\lambda_{2}\geq 0,$ with $\lambda_{3}=1-\lambda_{1}-\lambda_{2}>0,\beta\in(0,1),\delta\geq 0$, $\eta,L,U\in\mathbb{R}^{n}$ with $L<U$ element-wise, $d_{j}\in\mathbb{R}^{n}$ for each $j\in[m],A\succeq 0,$ and $\rho\geq 1$. Let $\left\{\left(x^{(j)},y^{(j)},z^{(j)},w^{(j)},\gamma^{(j)}\right)\right\}$ be a sequence generated by Algorithm 2 for solving ($P$). Then, the following holds:
@@
    Suppose that Robinson’s constraint qualification condition given in (18) holds at $(x^{*},\gamma^{*})$ with the index subset $\mathcal{L}$ indicated above. Then, $(x^{*},\gamma^{*})$ is a Lu–Zhang minimizer of ($P$).
@@
$\lambda_{1}{x^{(j)}}^{T}Ax^{(j)}-\lambda_{2}\Psi^{T}x^{(j)}+\lambda_{2}\delta\|z^{(j)}-\phi\|_{1}+\lambda_{3}(\gamma+\frac{1}{m(1-\beta)}\sum_{j=1}^{m}(-d_{j}^{T}w^{(j)}-\gamma)^{+})\geq\hat{C}$$\infty$.
@@
[…]
Since for each $j_{\ell}$ we have $y^{(j_{\ell})}_{i}\geq 0;\forall i\in\Gamma^{+}$ and $y^{(j_{\ell})}_{i}\leq 0;\forall i\in\Gamma^{-}$, from part (i), we can conclude $x^{(j_{\ell})}_{i}\geq 0;\forall i\in\Gamma^{+}$ and $x^{(j_{\ell})}_{i}\leq 0;\forall i\in\Gamma^{-}$. Further, $\textbf{e}^{T}x^{(j_{\ell})}=1$ and $x^{(j_{\ell})}-L\geq 0$ with $U-x^{(j_{\ell})}\geq 0$. Next, by injecting the second, third, and fourth equations of (21) into its first one, we obtain the following:
@@
Therefore, (22) reduces to:
$2\lambda_{1}Ax^{(j_{\ell})}-\lambda_{2}\Psi+t^{(j_{\ell})}\textbf{e}$ - $\underbrace{\begin{bmatrix}2\rho^{(j_{\ell})}{[x^{-}_{\Gamma^{+}};0]}^{(j_{\ell})}_{\mathcal{L}}\\ 0\end{bmatrix}}_{:=T_{5}^{(j_{\ell})}}+\underbrace{\begin{bmatrix}2\rho^{(j_{\ell})}{[x^{+}_{\Gamma^{-}};0]}^{(j_{\ell})}_{\mathcal{L}}\\ 0\end{bmatrix}}_{:=T_{6}^{(j_{\ell})}}+\underbrace{\begin{bmatrix}0\\ 2\rho^{(j_{\ell})}(x^{(j_{\ell})})_{\mathcal{L}^{c}}\end{bmatrix}}_{:=\Omega^{(j_{\ell})}}$ (23)
@@
We next show that $\{(t^{(j_{\ell})},T_{1}^{(j_{\ell})}+T_{3}^{(j_{\ell})},T_{2}^{(j_{\ell})}+T_{4}^{(j_{\ell})},T_{5}^{(j_{\ell})},T_{6}^{(j_{\ell})},\Omega^{(j_{\ell})}\}$ is bounded under Robinson’s condition on $(x^{*},\gamma^{*})$. Suppose not, and consider the following normalized sequence:
@@
[…]
$\widetilde{t}^{*}\textbf{e}-\widetilde{T}_{5}^{*}+\widetilde{T}_{6}^{*}+\widetilde{\Omega}^{*}-\widetilde{T}_{1}^{*}+\widetilde{T}_{2}^{*}-\widetilde{T}_{3}^{*}+\widetilde{T}_{4}^{*}\ni 0,$ (24)
[…]
$(\widetilde{t}^{*})^{2}-(\widetilde{T}^{*}_{5}-\hat{v})^{T}\widetilde{T}^{*}_{5}+(\tilde{v}-\widetilde{T}^{*}_{6})^{T}\widetilde{T}^{*}_{6}-\|\widetilde{\Omega}^{*}\|_{2}^{2}-(\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3}-v)^{T}(\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3})+(\bar{v}-\widetilde{T}^{*}_{2}-\widetilde{T}^{*}_{4})^{T}(\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4})=0,$
@@
$\|((\widetilde{t}^{*})^{2},\widetilde{T}^{*}_{5},\widetilde{T}^{*}_{6},\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3},\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4},\widetilde{\Omega}^{*})\|_{2}^{2}=\underbrace{\hat{v}^{T}\widetilde{T}_{5}^{*}}_{\leq 0}+\underbrace{\tilde{v}^{T}\widetilde{T}^{*}_{6}}_{\leq 0}+\underbrace{v^{T}(\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3})}_{\leq 0}+\underbrace{\bar{v}^{T}(\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4})}_{\leq 0};$
implying that $\|((\widetilde{t}^{*})^{2},\widetilde{T}^{*}_{5},\widetilde{T}^{*}_{6},\widetilde{T}^{*}_{1}+\widetilde{T}^{*}_{3},\widetilde{T}^{*}_{2}+\widetilde{T}^{*}_{4},\widetilde{\Omega}^{*})\|_{2}^{2}=0$. This is a contradiction! Consequently, the sequence $\{(t^{(j_{\ell})},T_{1}^{(j_{\ell})}+T_{3}^{(j_{\ell})},T_{2}^{(j_{\ell})}+T_{4}^{(j_{\ell})},T_{5}^{(j_{\ell})},T_{6}^{(j_{\ell})},\Omega^{(j_{\ell})}\}$ is bounded and has a convergent subsequence. Let $\{(\bar{t},\bar{T}_{1},\bar{T}_{2},\bar{T}_{5},\bar{T}_{6},\bar{\Omega})$ be the limit point of this sequence. By passing the limit in (23) and applying the results from part (i), one can see that the KKT conditions (20) are satisfied.
So far, we have proved that all the KKT conditions in (20) hold at $(x^{*},\gamma^{*})$, except for its second equation. To show this, recall the fourth equation in (21):
$\lambda_{3}-\frac{1}{m(1-\beta)}\sum_{j=1}^{m}\partial\left(.)^{+}\right|_{-d_{j}^{T}w^{(j_{\ell})}-\gamma^{(j_{\ell})}}\ni 0.$
By taking the limit as $j_{\ell}\to\infty$, the boundedness of $\gamma^{(j_{\ell})}$ shown in (3.1), the continuity of $\max(x,0)$, and considering part (i), we see that
$\lambda_{3}-\frac{1}{m(1-\beta)}\sum_{j=1}^{m}\partial\left(.)^{+}\right|_{-d_{j}^{T}x^{*}-\gamma^{*}}\ni 0.$
Therefore, we showed that all the KKT conditions hold at $(x^{*},\gamma^{*})$.
Under Theorem 3.34 of [15] and Theorem 2.3 of [12], by virtue of the linearity of constraints in ($P$), except for the cardinality constraint, we conclude that Algorithm 2 converges to a Lu–Zhang minimizer of this problem under the given assumptions. ∎